A 20 mm diameter, 500 g cylinder-shaped drill is spun up clockwise, starting from rest, by an applied torque of -0.2 [N m], after 2 seconds.

a. What is the drill's angular speed?

b. Through what total angle has it spun?
Answer #1

(a) moment of inertia of cylinder = 1/2mr^2 = 0.5*0.5kg*(0.01 m)^2 = 0.000025 kgm2
torque = Ia
a = angular acceleration = 0.2/0.000025 = 8000 rad/s2
w = at = 8000*2 = 16000 rad/s
(b) angle theta = 1/2at^2 = 0.5*8000*4 = 16000 radians